The vertices of a quadrilateral are (-1, -1), (1,7), (6,5 ), and (3, -1). What is the area?
Answer by Alan3354(69443)

The vertices of a quadrilateral are (-1, -1), (1,7), (6,5 ), and (3, -1). What is the area?
--------------
Label them.
The vertices of a quadrilateral are A(-1,-1), B(1,7), C(6,5), and D(3,-1). What is the area?
------
A. B. C. D. A
-1 +1 +6 +3 -1
-1 +7 +5 -1 -1
---------
Add the diagonal products starting at the upper left = -7 + 5 - 6 - 3 = -11
----
Add the diagonal products starting at the lower left = -1 + 42 + 15 + 1 = 57
---
Area = 1/2 the difference = 1/2 * 68
= 34 sq units